OpenStudy (anonymous):

Find the Inverse Laplace transform of s^2/((s^2+a^2)(s^2+b^2)).

OpenStudy (anonymous):

Decompose into partial fractions: \[\frac{s^2}{(s^2+a^2)(s^2+b^2)}=\frac{Cs+D}{s^2+a^2}+\frac{Es+F}{s^2+b^2}\] (C,D,E,F are used to avoid confusion with a,b) \[s^2=(Cs+D)(s^2+b^2)+(Es+F)(s^2+a^2)\\ s^2=(C+E)s^3+(D+F)s^2+(Cb^2+Ea^2)s+(Db^2+Fa^2)\] yielding the system \[\begin{cases}C+E=0\\D+F=1\\Cb^2+Ea^2=0\\Db^2+Fa^2=0\end{cases}\] Need help solving for the constants @cupcake1123 ?

OpenStudy (anonymous):

I'll try to solve for the constants, and then will check with you.. Thanks so much!

OpenStudy (anonymous):

can we solve this by using the table?

OpenStudy (anonymous):

can you help me with the constants please?

OpenStudy (anonymous):

From the first equation, you have \(-C=E\). Plug this into the third equation: \[Cb^2+(-C)a^2=0\\ C(b^2-a^2)=0\] Assuming \(a\not=b\), you have \(C=0\), and in turn, \(E=0\). (Is this the right assumption? I would think so; otherwise, there'd be no reason to differentiate between the two.) So you're left with solving the following system: \[\begin{cases}D+F=1\\Db^2+Fa^2=0\end{cases}\] Multiply the first equation by \(-a^2\): \[\begin{cases}-Da^2-Fa^2=-a^2\\Db^2+Fa^2=0\end{cases}\] Adding the equation together gives you \[Db^2-Da^2=-a^2\\ D(b^2-a^2)=-a^2\\ D=\frac{a^2}{a^2-b^2}\] So, you also get \(F=1-\dfrac{a^2}{a^2-b^2}\). So partial fraction decomp yields \[\frac{a^2}{a^2-b^2}\cdot\frac{1}{s^2+a^2}+\left(1-\frac{a^2}{a^2-b^2}\right)\cdot \frac{1}{s^2+b^2}\] Can you take the inverse transform now?

OpenStudy (anonymous):

do i solve that using the table?

OpenStudy (anonymous):

You could, sure. \[\mathcal{L}^{-1}\left\{\frac{1}{s^2+a^2}\right\}=\cdots\] There's some algebraic manipulation you'll have to do first in order to make a proper transformation.

OpenStudy (anonymous):

okay.. but is there another approach.. i need to know two ways.. i know the one using the table.

OpenStudy (anonymous):

I think there's an actual integral formula for the inverse transform, but I've only ever had to refer to a table of transforms to do these kinds of problems

OpenStudy (anonymous):

okay.. i have a question.. can we take individual ones in the multiplication and assign the transform table values.. for instance, if one has laplace inverse of e^at and the one being mutliplied to it has cos(at), can we mutliply the inverses?

OpenStudy (anonymous):

Do you mean \[\mathcal{L}^{-1}\left\{e^{at}\right\}\cdot\mathcal{L}^{-1}\left\{\cos(at)\right\}=\mathcal{L}^{-1}\left\{e^{at}\cos(at)\right\}~~?\] No, the Laplace transform and its inverse don't work that way with multiplication.

OpenStudy (anonymous):

If you meant something else, could you draw it? I'm not sure what you mean.

OpenStudy (anonymous):

yes.. that was my question.. so how does it work?

OpenStudy (anonymous):

The property you're thinking of only works with addition and scalar multipliers: \[\mathcal{L}^{-1}\left\{cF(s)+dG(s)\right\}=c\mathcal{L}^{-1}\{F(s)\}+d\mathcal{L}^{-1}\{G(s)\}\] This means the Laplace transform and its inverse are "linear" operators. You can derive this property by using the integral definition of the transform.

OpenStudy (anonymous):

Just to address your example: \[\mathcal{L}\left\{e^{at}\cos(bt)\right\}=\frac{s-a}{(s-a)^2+b^2}\] (I should have removed the "inverse" notation before.) So basically, if you're given a function multiplied by \(e^{at}\), \[f(t)=e^{at}g(t)\], then the transform is the transform of the function, shifted by \(a\): \[\mathcal{L}\{e^{at}g(t)\}=G(s-a)\] (where \(G(s)\) is the Laplace tranform of \(g(t)\).)